Charles company used the percent of sales method to determine its bad debts expense. At the end of the current year, the company's unadjusted trial balance reported the following selected amounts:

Accounts Receivable                              $245,000 debit

Allowance for doubtful Accounts         $300 credit

Net Credit Sales                                     $900,000

Based on past experience, the company estimates 0.5% of credit sales to be uncollectible. What amount would be reported in the Balance Sheet under Accounts Receivable, net of doubtful accounts?

A) $245,000

B) $244,700

C) $240,200

D) $240,500

E) $240,800
Answer #1

Bad debt expense = 900000*.50% = $4500

Allowance for doubtful accounts after adjustment = 4500+300 = $4800

Balance of account receivable, net of doubtful accounts = 245000-4800 = $240200

So answer is c) $240200